OpenStudy (anonymous):

Evaluate 48 + 24 ÷ 8 - 22

OpenStudy (anonymous):

29 is the answer

OpenStudy (texaschic101):

48 + 24/8 - 22 = 48 + 3 - 22 = 51 - 22 = 29

OpenStudy (anonymous):

there is no 29 .-. like its not any of the options

OpenStudy (texaschic101):

Is there any parenthesis in your original problem ??

OpenStudy (anonymous):

no :(

OpenStudy (texaschic101):

I don't know what to say...29 is the answer. You would do division first, then do add/subtract from left to right.